OpenStudy (anonymous):

f(x)=ln(sinx) find f'(pi/4) a. -.5 ln2 b. (sqrt2)/2 c. 0 d. 1 e. undefined please help

OpenStudy (anonymous):

f(x)=ln(sinx) find f'(pi/4) let u=sin x f '(x)=(1/sin x)(cos x) =cos x/sin x=cot x =1/tan x, f'(pi/4) =1/tan(pi/4)=1